What is thermal resistance of a conductor? On factors does it depend?
 

Grade:10

6 Answers

mahendra
8 years ago
Thermal resistance (R):
The thermal resistance of a body is a measure of its opposition to the flow of heat through it.
It is defined as the ratio of temperature difference to the heat current.
R=L/KA
Thermal resistance of a body depends on length and area of cross-section of the conductor.
 
salvathshaik
8 years ago
The thermal conductivity of a material may depend on temperature. ... thermal resistance and the reciprocal property, thermal conductancecan be calculated. .... is less thermally conductive than diamond, which is an electrical insulator, but due ... by a complexity factor CF (defined as the number of atoms/primitive unit cell).
 
mahendra
8 years ago
 
Thermal resistance (R):
The thermal resistance of a body is a measure of its opposition to the flow of heat through it.
It is defined as the ratio of temperature difference to the heat current.
R=L/KA
Thermal resistance of a body depends on length and area of cross-section of the conductor.
 
mahendra
8 years ago
 
Thermal resistance (R):
The thermal resistance of a body is a measure of its opposition to the flow of heat through it.
It is defined as the ratio of temperature difference to the heat current.
R=L/KA
Thermal resistance of a body depends on length and area of cross-section of the conductor.
 
mahendra
8 years ago
 
Thermal resistance (R):
The thermal resistance of a body is a measure of its opposition to the flow of heat through it.
It is defined as the ratio of temperature difference to the heat current.
R=L/KA
Thermal resistance of a body depends on length and area of cross-section of the conductor.
 
mahendra
8 years ago
 
Thermal resistance (R):
The thermal resistance of a body is a measure of its opposition to the flow of heat through it.
It is defined as the ratio of temperature difference to the heat current.
R=L/KA
Thermal resistance of a body depends on length and area of cross-section of the conductor.